OpenStudy (anonymous):

if a^2-b^2=21and a^2+b^2=29,which of the value of ab

OpenStudy (anonymous):

a=25, b=2

OpenStudy (anonymous):

wuhuuuuuuuuuuuuu :D

OpenStudy (anonymous):

a^2-b^2=21 a^2+b^2=29 adding we get,2a^2=50 a^2=25 a=5 similarly,subtracting,-2b^2=-8 b^2=4 b=2 ab=5*2 ab=10

OpenStudy (anonymous):

Strictly speaking, plus minus 2 and 5.